Grade 11Mechanics

kinetic energy of a block of mass 2 kg is given as kinetic energy 16t^4 and the block is moving in a circular path of radius 8 m.find the power generated by the applied force at t=0.5 sec

Raghul
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
Since power=force×velocityAcc/q Ke=16t^4 i.e. mv^2=32t^4i.e. At t=1/2 mv^2=2And Force=mv^2/ri.e. F=2/8Since, mv^2 =2 =>v^2=1 (since m=2)so, v=1Now, power=force× velocity=>P=2/8×1=2/8 watt
 
Regards
Arun